22. If f(x)=x^2 and g(x)=x+2, find f[g(x)]
23. If f(x)=x^3+4 and g(x)=x+3, find [g•f](2)
24. Find the inverse of the function f(x)=4x+1.
25. Which of the following pairs of functions are inverse functions?
a. f(x)=3x-9, g(x)=-3x+9
b. f(x)=4-x, g(x)=4+x
c. f(x)=x+5, g(x)=x-5
d. f(x)=x, g(x)= -x



Answer by khwang(438)

22. f[g(x)] = f(x+2) = (x+2)^2.
23. [g•f](2)= g(f(2)) = g(2^3 + 4) = g(12) = 12+3 = 15
24. Let y = f(x)=4x+1, then 4x = y - 1, or x =(y - 1)/4.
Exchange x,y we have y =(x - 1)/4.
Hence, the inverse function is g(x) = (x - 1)/4.
25. Answer : c. f(x)=x+5, g(x)=x-5
Since f[g(x)] = g[f(x)] = x